Solve using any method! y=-0.7x+6.9
y=1.2x-6.4

Answer by checkley77(12844)

y=-0.7x+6.9
-(y=1.2x-6.4) subtract
------------------
0=-1.9x+13.3
1.9x=13.3
x=13.3/1.9
x=7 ans.
y=-.7*7+6.9
y=-4.9+6.9
y=2 ans.
Proof:
2=1.2*7-6.4
2=8.4-6.4
2=2